After graduation, you and 44 of your closest friends (assume 22 males and 22 females) charter a plane to Hawaii. Unfortunately, you all crash land (safely) ona deserted island somewhere off the coast of the US. Unfortunately, no one finds you and you start a new population totally isolated from the rest of the world. Unbeknownst to them, three of your friends carry (are heterozygotes) for a recessive allele (g) that is lethal in later stages of life (past the reproductive stage). Assuming that the frequency of this allele does not change as the population grows, what will be the expected frequency of this lethal, recessive disease on your island? Show your work.
Answer #1

3 friends out of 44 are heterozygous for the lethal allele. The genotypes of heterozygotes is Gg. According to Hardy-Weinberg's law, genotype frequency of heterozygous population = 2pq = 3/44 = 6.8%. So, genotype frequency of dominant population,GG= p ^ 2 = 93.2%.

Frequency of dominant allele = p = under root (93.2) = 9.65

Frequency of recessive allele = q = 1-p = 1 - 9.65 = 0.35

Frequency of this lethal disease = q^2 = 0.35^2 = 0.1225 = 0.12% ~ 0.1%